# HistoPrism: Unlocking Functional Pathway Analysis from Pan-Cancer Histology via Gene Expression Prediction

### v1

@@ 3.1 Problem Formulation
We consider an H&E-stained whole-slide image, divided into $N$ non-overlapping patches. Each patch is represented by a feature vector $\mathbf{x}_{i}\in\mathbb{R}^{D_{img}}$, extracted by a pre-trained pathology foundation model (PFM). Spatial transcriptomics provides the corresponding raw count vector of gene expressions, which we normalize as $\mathbf{y}_{i}\in\mathbb{R}^{D{gene}}$ using a log1p transformation. Additionally, each slide is associated with a global condition cancer type, encoded as a one-hot vector $\mathbf{c}\in\{0,1\}^{D_{onco}}$.
@@ 3.2 HistoPrism: A Direct-Mapping Architecture
HistoPrism is trained end-to-end by minimizing the Mean Squared Error (MSE) $\mathcal{L}_{\text{MSE}}$ between the predicted and ground-truth gene expression values across all $N$ patches.
@@ 3.3 A Framework for Evaluating Biological Coherence
The standard protocol in this domain is to evaluate the Pearson Correlation Coefficient (PCC) between predicted and ground-truth expression for the top-$N$ most highly-variant genes (HVGs) across a test set. While this metric is useful for gauging a model’s ability to capture high-magnitude signals, its clinical and biological relevance is limited. It focuses on a small, statistically-driven subset of genes, ignoring thousands of others, and it fails to measure whether a model has learned the coordinated expression patterns that define a functional biological process. A model can achieve a high HVG while failing to generate biologically coherent predictions, thus limiting its translational potential.
@@
    Source Curation: We begin by aggregating a comprehensive set of pathways from two authoritative, widely-used biological databases: the Hallmark gene sets from the Molecular Signatures Database (MSigDB) (Broad Institute, 2025), which represent well-defined biological states or processes, and the Gene Ontology (GO) database (Gene Ontology Consortium, 2025), from which we include terms for Biological Process (BP), Cellular Component (CC), and Molecular Function (MF).
@@
For each pathway, the evaluation score is the computed across all of its member genes. Let $\mathcal{Y}=\{(\mathbf{y}_{i},\hat{\mathbf{y}}_{i})\}_{i=1}^{N}$ denote the paired ground-truth and predicted gene expression sets for $N$ whole-slide images (WSIs). Each WSI $i$ contains $n_{i}$ patches, with $\mathbf{y}_{i}=[\mathbf{y}_{i1},\dots,\mathbf{y}_{in_{i}}]^{\top}$ and $\hat{\mathbf{y}}_{i}=[\hat{\mathbf{y}}_{i1},\dots,\hat{\mathbf{y}}_{in_{i}}]^{\top}$, where $\mathbf{y}_{ij},\hat{\mathbf{y}}_{ij}\in\mathbb{R}^{D_{\text{gene}}}$ represent the expression vectors of $D_{\text{gene}}$ genes.
@@
where $\mathbf{y}_{i,:,g}=[y_{i1g},\dots,y_{in_{i}g}]^{\top}$ and $\hat{\mathbf{y}}_{i,:,g}=[\hat{y}_{i1g},\dots,\hat{y}_{in_{i}g}]^{\top}$ denote the expression profiles of gene $g$ across all patches of WSI $i$.
@@ 4.1 Experimental Setup
Our proposed model HistoPrism consists of $1$ cross attention layer with $4$ heads and $2$ transformer layers with $8$ heads and $256$ hidden dimension receptively. HistoPrism is trained end-to-end with UNI PFM (Chen et al., 2024b) with a gene panel of size 38,982 curated by STPath. The training details are included in Appendix  B.

### v2

@@ 3.1 Problem Formulation
We consider an H&E-stained whole-slide image, divided into NN non-overlapping patches. Each patch is represented by a feature vector $\mathbf{x}_{i}\in\mathbb{R}^{D_{img}}$, extracted by a pre-trained pathology foundation model (PFM). Spatial transcriptomics provides the corresponding raw count vector of gene expressions, which we normalize as $\mathbf{y}_{i}\in\mathbb{R}^{D{gene}}$ using a log1p transformation. Additionally, each slide is associated with a global condition cancer type, encoded as a one-hot vector $\mathbf{c}\in\{0,1\}^{D_{onco}}$.
@@ 3.2 HistoPrism: A Direct-Mapping Architecture
HistoPrism is trained end-to-end by minimizing the Mean Squared Error (MSE) $\mathcal{L}_{\text{MSE}}$ between the predicted and ground-truth gene expression values across all NN patches.
@@ 3.3 A Framework for Evaluating Biological Coherence
The standard protocol in this domain is to evaluate the Pearson Correlation Coefficient (PCC) between predicted and ground-truth expression for the top-NN most highly-variant genes (HVGs) across a test set. While this metric is useful for gauging a model’s ability to capture high-magnitude signals, its clinical and biological relevance is limited. It focuses on a small, statistically-driven subset of genes, ignoring thousands of others, and it fails to measure whether a model has learned the coordinated expression patterns that define a functional biological process. A model can achieve a high HVG while failing to generate biologically coherent predictions, thus limiting its translational potential.
@@
    Source Curation: We begin by aggregating a comprehensive set of pathways from two authoritative, widely-used biological databases: the Hallmark gene sets from the Molecular Signatures Database (MSigDB) (Broad Institute, 2025), which represent well-defined biological states or processes, and the Gene Ontology (GO) database (Gene Ontology Consortium, 2025), from which we include terms for Biological Process (BP), Cellular Component ($C$), and Molecular Function (MF).
@@
For each pathway, the evaluation score is the computed across all of its member genes. Let $\mathcal{Y}=\{(\mathbf{y}_{i},\hat{\mathbf{y}}_{i})\}_{i=1}^{N}$ denote the paired ground-truth and predicted gene expression sets for NN whole-slide images (WSIs). Each WSI $i$ contains $n_{i}$ patches, with $\mathbf{y}_{i}=[\mathbf{y}_{i1},\dots,\mathbf{y}_{in_{i}}]^{\top}$ and $\hat{\mathbf{y}}_{i}=[\hat{\mathbf{y}}_{i1},\dots,\hat{\mathbf{y}}_{in_{i}}]^{\top}$, where $\mathbf{y}_{ij},\hat{\mathbf{y}}_{ij}\in\mathbb{R}^{D_{\text{gene}}}$ represent the expression vectors of $D_{\text{gene}}$ genes.
@@
where $\mathbf{y}_{i,:,g}=[y_{i1g},\dots,y_{in_{i}g}]^{\top}$ and $\hat{\mathbf{y}}_{i,:,g}=[\hat{y}_{i1g},\dots,\hat{y}_{in_{i}g}]^{\top}$ denote the expression profiles of gene $g$ across all patches of WSI ii.
@@ 4.1 Experimental Setup
Our proposed model HistoPrism consists of 11 cross attention layer with 44 heads and 22 transformer layers with 88 heads and 256256 hidden dimension receptively. HistoPrism is trained end-to-end with UNI PFM (Chen et al., 2024b) with a gene panel of size 38,982 curated by STPath. The training details are included in Appendix  B.